# MultiHal: Multilingual Dataset for Knowledge-Graph Grounded Evaluation of LLM Hallucinations

### v1

@@ Abstract
Large Language Models (LLMs) have inherent limitations of faithfulness and factuality, commonly referred to as hallucinations. Several benchmarks have been developed that provide a test bed for factuality evaluation within the context of English-centric datasets, while relying on supplementary informative context like web links or text passages but ignoring the available structured factual resources. To this end, Knowledge Graphs (KGs) have been identified as a useful aid for hallucination mitigation, as they provide a structured way to represent the facts about entities and their relations with minimal linguistic overhead. We bridge the lack of KG paths and multilinguality for factual language modeling within the existing hallucination evaluation benchmarks and propose a KG-based multilingual, multihop benchmark called MultiHal framed for generative text evaluation. As part of our data collection pipeline, we mined 140k KG-paths from open-domain KGs, from which we pruned noisy KG-paths, curating a high-quality subset of 25.9k. Our baseline evaluation shows an absolute scale increase by approximately 0.12 to 0.36 points for the semantic similarity score in KG-RAG over vanilla QA across multiple languages and multiple models, demonstrating the potential of KG integration. We anticipate MultiHal will foster future research towards several graph-based hallucination mitigation and fact-checking tasks.
[Uncaptioned image] Code: https://github.com/ernlavr/multihal
[Uncaptioned image] Data: https://huggingface.co/datasets/ernlavr/multihal
@@ 1 Introduction
Factual inconsistencies in LLM outputs, commonly referred to as hallucinations, are often a bottleneck for production-grade deployment of LLM systems hallucSurvey . Although hallucinations may be beneficial for tasks involving creativity jiang2024survey or even drug discovery yuan2025hallucinations , they become a liability for other tasks that require factually consistent outputs, for example, information retrieval, summarization and question answering lavrinovics . Additionally, hallucSurvey ; Augenstein2024 suggests that hallucinations impair the trust and usefulness of AI systems, and even pose certain societal risks by enabling the generation of convincing misinformation Augenstein2024 ; rogers_news_farms . Hallucinations can stem from multiple shortcomings in model training, such as reinforcement learning from human feedback (RLHF) bai2022training , in cases when human preferences are towards non-factual answers zhang2023siren , instruction tuning where given instructions exceed a model’s knowledge boundary zhang2023siren ; halluc_survey , or due to lack of up-to-date knowledge. Furthermore, hallucinations occur with varied levels of frequency and intensity depending on the generated language chataigner2024multilingual ; qi-etal-2023-cross . A general trend is observed that, in terms of factual consistency, English outputs are the most stable and overall factual quality decreases with lower resourced languages. This varied degree of factuality across languages only further impairs the usability and inclusiveness of LLMs in different applications.
@@
To this end, Retrieval Augmented Generation (RAG) ragtruth ; zhao2024retrieval is the most widely adopted method for improving factuality, which supplements the input query to an LLM with relevant text passages to improve the factuality of LLM outputs. The main advantage of RAG is that it does not require retraining the generator LLM, a process that is time-consuming and resource-intensive. However, RAG is still limited by the LLM context window size zhao2024retrieval , its sensitivity to input prompt formatting whatSota ; maia2024efficient , and Needle in a Haystack problem gao2025u , where important details can be lost in a large pool of text.
@@
KG-RAG sanmartin2024kg ; peng2024graph provides several advantages over document-based RAG and has also been suggested as a promising methodology for limiting LLM hallucinations pan2024unifying ; russaKgSurvey , primarily leveraging the structural and factual qualities of the KGs through sets of KG-paths that describe entities and their relationships with minimal linguistic overheads. Furthermore, KG integration within language modeling can alleviate the need for full re-training when utilized during inference sun2023think ; luo2024rog or post-generation guan2024mitigating . This is valuable for use-cases with rapidly developing knowledge or limited computational resourceslavrinovics . The structured, factually rich and linguistically minimal nature of the KGs can potentially decrease the risks of the Needle-in-a-Haystack problem and limitations of the context window size. Conditioning LLMs on KGs can also enable optimal output scrutiny and explainability by allowing the outputs to be traced back to explicit sources, making cross-checking less time-consuming than document-based RAG. Furthermore, KGs accompany each entity with rich metadata, but their optimal use in factual language modeling is still an open question.
@@
Although KG-RAG is rapidly gaining attention to improve the factuality in LLM, existing QA benchmark data sets zhao2023felm ; lin2021truthfulqa ; li2023halueval ; simpleQA ; rahman2024defan ; shroom2024 ; halubench on LLM hallucinations rely primarily on textual data for contextual information and provide no multilingual support. While the questions in these benchmark datasets are compiled from different sources, the answers for FELM zhao2023felm HaluEval  li2023halueval Shroom2024  shroom2024 are LLM-generated and evaluated using LLM-as-a-judge or human annotation. For some datasets lin2021truthfulqa ; zhao2023felm ; simpleQA , the answers are supported with external contextual information from textual resources such as webpages. Therefore, in this paper, we bridge these critical gaps by presenting a novel multilingual hallucination benchmark MultiHal, grounded on factual information from Wikidata vrandevcic2014wikidata KG. MultiHal is based on a total of 7 common benchmarks that lack structured factual and multilingual coverage, namely Felm zhao2023felm , TruthfulQAlin2021truthfulqa (TQA), HaluEvalli2023halueval , HaluBench halubench , SimpleQA simpleQA , DefAn rahman2024defan , Shroom2024 shroom2024 . We propose a data collection framework as illustrated in Figure 1, to aggregate over 31k unique questions from aforementioned datasets, enriching them by mining 140k KG paths and ensuring factual consistency by filtering using LLM-as-a-judge. To enable multilingual hallucination evaluation, our compiled dataset comprising questions, ground-truth answers and KG paths are translated to Spanish, French, Italian, Portuguese and German. Therefore, our main contributions are as follows:
@@ 2 MultiHal
MultiHal builds upon a set of 7 previously established benchmarks by enriching them with factual information in the form of relevant paths from Wikidata. The choice of these benchmarks is motivated by their relevancy to factuality evaluation, yet they lack support for factual grounding of the answers, leveraging KG and LLM integration models, and multilingual evaluation. We summarize the basic dataset statistics in Table 1, for a dataset schema description see Appendix A. These foundational benchmarks are all filtered for generative question-answering based on general/trivia domains, and a full breakdown of domains for our final dataset with the fine-grained domains. Furthermore, datasets such as Shroom2024 shroom2024 , FELM zhao2023felm , HaluEval li2023halueval , HaluBench halubench are primarily oriented towards evaluating hallucination detection models consisting of both hallucinated and non-hallucinated data, therefore, the data is repurposed by filtering for rows labelled as non-hallucinated. We consider that each unique question-path pair is a data point. The count difference between data points and unique questions is because we mine multiple candidate paths per question. The overview of each of the processing stages in our dataset collection pipeline is illustrated in Figure 1. The following sections scope into the methodological details of each of the processing stages of the proposed dataset collection framework. Additionally we report on our computing processing times and CO2 emissions in Appendix J. Our original contributions are released under CC-BY-4.0 license terms.
@@
Shroom2024 Definition Modeling CC-BY 346 160 1 170.86 73.37
Felm World knowledge CC-BY-NC-SA-4.0 73 17 1 95.25 75.26
MultiHal (total) - CC-BY-4.0 25,905 7095 48 106.27 70.98
@@ 2.1 Dataset Preprocessing
Considering that MultiHal builds upon established benchmarks, question deduplication is performed to avoid data leakage across the foundations. Deduplication is based on computing sentence embeddings using SentenceTransformers [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 reimers-2019-sentence-bert and computing all possible pair-wise cosine similarity between the questions. The ground-truth answers and any present supplementary context of the pair of questions with a sentence similarity threshold above 0.99 are merged. Deduplication was exclusively skipped for DefAn-QSRanking subset due to a large amount of questions consisting of nearby years for corresponding university rankings, which led to a very high number of false positives among the data points.
Additionally, we discard data points where the ground-truth answers are phrases such as "I have no comment", which indicate refusal to answer, and we define them as refusal types. We compile a list of refusals consisting of a list of text patterns as described in B. Any rows with output columns that exactly match, case-insensitively, one of these refusal phrases are filtered out.
@@ 2.2 KG Path Mining
The overall idea is to mine relevant paths from Wikidata vrandevcic2014wikidata by extracting the core semantic entities from a given question $\mathcal{Q}$ and its ground-truth answer $\mathcal{A}$, and afterwards matching them to Wikidata entities. Afterwards, Wikidata is queried for existing paths between the extracted entities in $\mathcal{Q}$ and $\mathcal{A}$.
Entity Matching from Text to Knowledge Graphs. The core entity extraction and matching from raw text is based on Falcon 2.0 falcon2 . The authors of Falcon 2.0 make their processing engine available via an API [^2] https://labs.tib.eu/falcon/falcon2/ which we call to retrieve subjects from question $\mathcal{Q}$ and objects from answer $\mathcal{A}$. Given a text passage, Falcon 2.0 outputs a ranked list of entities as candidates in Wikidata, we use the Top-3 candidates. Additionally, we noted that not always Falcon 2.0 retrieves relevant Wikidata entities although it matches them correctly to DBpedia. Therefore use Falcon 2.0 to additionally return DBpedia entities, which we then map back to Wikidata using the query in Listing 1 in Appendix C.
Additionally, foundational benchmarks such as FELM zhao2023felm , SimpleQA simpleQA , TruthfulQA lin2021truthfulqa contain supplementary context in the form of Wikipedia links which we map to Wikidatavrandevcic2014wikidata entities using Wikipedia public API [^3] https://en.wikipedia.org/w/api.php?action=query&prop=pageprops&titles=\$WIKIPEDIA_ID&format=json. The Wikipedia-to-Wikidata retrieval is done by taking the page title embedded in the given Wikipedia link and replacing it with the \$WIKIPEDIA_ID placeholder.
@@
Knowledge Graph Querying. We query Wikidata in order to find existing paths between the extracted subject-object entities up to 2 hops. As additional pre-processing steps before querying, we remove circular subject-object pairs where the subject-object is identical, as well as inversion of subject-object to accommodate for the directionality of the Wikidata graph.
Depending on the foundational benchmark, we create custom queries for the different answer types we encounter when merging all our foundational benchmarks. The answer type is is denoted by answer_type column in MultiHal, see the schema in Appendix A Table 6. In Appendix C, see Listing 2 for Wikidata entity query, Listing 3 for date-literal query and Listing 4 for numerical-literal query.
The answer types, such as numericals and dates, are queried with value limitations for numerical and time-based properties in the final hop, as shown in Listings 4 and 3 to improve query speed. See Appendix D Listing 6 for time-based properties and Listing 7 for numerical properties. For querying, we use the public Wikidata endpoint [^4] https://query.wikidata.org/bigdata/namespace/wdq/sparql, our path cut-off date is April 2025.
For decoding the Wikidata entity labels, we run a separate pass using the query in Appendix C Listing 5. When querying for labels, we discard any statements, entities or objects that cannot be directly mapped to natural language text labels.
@@ 2.3 KG Path Quality Evaluation: LLM as Judge
As a method for filtering out noisy KG paths and identifying high-quality paths, we employ a two-step LLM-as-a-judge methodology li2025generationjudgmentopportunitieschallenges ; yu2024rankrag : firstly, for questions with more than 10 candidate paths the top-10 paths selection is done to limit the total count; secondly, scoring each path individually to identify low and high quality paths. For both selection and scoring we use GPT-4o Mini from OpenRouter API [^5] https://openrouter.ai/openai/gpt-4o-mini with sampling temperature 0.1. The goal of the selection is to decrease the overall number of KG paths, resulting in a decrease from 140k to 25.9k paths.
Selection Step.We construct a prompt for selecting the 10 most relevant paths with respect to the question-answer, and available optional answer pairs. Selection is intended to be done without any particular ordering, see Listing 8 in Appendix E. The set of paths for each question is processed in two passes, and in both scenarios, the order of the paths is shuffled to avoid any ordering biases li2025generationjudgmentopportunitieschallenges . From the two passes, we consider only the overlapping paths from the two candidate sets as the final collection of paths.
During the selection phase, LLM-generated outputs are validated by checking for exact matches against the set of paths corresponding to the question. Any generated paths that do not match exactly are discarded to mitigate the risk of syntactic errors or hallucinations introduced by the LLM-as-a-judge. This process is repeated up to three times or until a total of 10 valid paths are obtained. If, after three attempts, fewer than 10 valid paths are selected, the remaining slots are filled by randomly sampling from the original KG path pool for the corresponding question. For questions that already contain 10 or fewer paths, the selection procedure is bypassed entirely.
Scoring Step.Once a set of candidate paths is established, we construct another prompt for rating their relevance with respect to the given question and answer, and we process each path individually, see Appendix E Listing 9 for the instructions. Scoring is done by determining the quality score on a scale of 1-5, where 1 indicates a path which is completely unrelated to the question and answer, and 5 indicates an explicit answer to the question. From our final benchmark we filter out all paths rated 1-3, which we deem as low-quality and leave only paths rated 4-5 as high-quality ones.

### v2

# MultiLingual Dataset for Knowledge-Graph Grounded Evaluation of LLM Hallucinations

@@ Abstract
Large Language Models (LLMs) have inherent limitations of faithfulness and factuality, commonly referred to as hallucinations. Several benchmarks have been developed that provide a test bed for factuality evaluation within the context of English-centric datasets, while relying on supplementary informative context like web links or text passages but ignoring the available structured factual resources. To this end, Knowledge Graphs (KGs) have been identified as a useful aid for hallucination mitigation, as they provide a structured way to represent the facts about entities and their relations with minimal linguistic overhead. We bridge the lack of KG paths and multilinguality for factual language modeling within the existing hallucination evaluation benchmarks and propose a KG-based multilingual, multihop benchmark called MultiHal framed for generative text evaluation. As part of our data collection pipeline, we mined 140k KG-paths from open-domain KGs, from which we pruned noisy KG-paths, curating a high-quality subset of 25.9k. Our baseline evaluation shows an absolute scale improvement by approximately 0.12 to 0.36 points for the semantic similarity score, 0.16 to 0.36 for NLI entailment and 0.29 to 0.42 for hallucination detection in KG-RAG over vanilla QA across multiple languages and multiple models, demonstrating the potential of KG integration. We anticipate MultiHal will foster future research towards several graph-based hallucination mitigation and fact-checking tasks.
Code:https://github.com/ernlavr/multihal
Data:https://huggingface.co/datasets/ernlavr/multihal
@@ 1 Introduction
Factual inconsistencies in LLM outputs, commonly referred to as hallucinations, are often a bottleneck for production-grade deployment of LLM systems (Huang et al., 2025a). Although hallucinations may be beneficial for tasks involving creativity (Jiang et al., 2024) or even drug discovery (Yuan & Färber, 2025), they become a liability for other tasks that require factually consistent outputs, for example, information retrieval, summarization and question answering (Lavrinovics et al., 2025). Additionally, Huang et al. (2025a); Augenstein et al. (2024) suggests that hallucinations impair the trust and usefulness of AI systems, and even pose certain societal risks by enabling the generation of convincing misinformation (Augenstein et al., 2024; Puccetti et al., 2024). Hallucinations can stem from multiple shortcomings in model training, such as reinforcement learning from human feedback (RLHF) (Bai et al., 2022), in cases when human preferences are towards non-factual answers (Zhang et al., 2023), instruction tuning where given instructions exceed a model’s knowledge boundary (Zhang et al., 2023; Huang et al., 2025b), or due to lack of up-to-date knowledge. Furthermore, hallucinations occur with varied levels of frequency and intensity depending on the generated language (Chataigner et al., 2024; Qi et al., 2023). A general trend is observed that, in terms of factual consistency, English outputs are the most stable and overall factual quality decreases with lower resourced languages. This varied degree of factuality across languages only further impairs the usability and inclusiveness of LLMs in different applications.
@@
To this end, Retrieval Augmented Generation (RAG) (Niu et al., 2024; Zhao et al., 2024) is the most widely adopted method for improving factuality, which supplements the input query to an LLM with relevant text passages to improve the factuality of LLM outputs. The main advantage of RAG is that it does not require retraining the generator LLM, a process that is time-consuming and resource-intensive. However, RAG is still limited by the LLM context window size (Zhao et al., 2024), its sensitivity to input prompt formatting (Mizrahi et al., 2024; Maia Polo et al., 2024), and Needle in a Haystack problem (Gao et al., 2025), where important details can be lost in a large pool of text.
@@
KG-RAG (Sanmartín, 2024; Peng et al., 2024) provides several advantages over document-based RAG and has also been suggested as a promising methodology for limiting LLM hallucinations (Pan et al., 2024; 2023), primarily leveraging the structural and factual qualities of the KGs through sets of KG-paths that describe entities and their relationships with minimal linguistic overheads. Furthermore, KG integration within language modeling can alleviate the need for full re-training when utilized during inference (Sun et al., 2023; Luo et al., 2024) or post-generation (Guan et al., 2024). This is valuable for use-cases with rapidly developing knowledge or limited computational resources (Lavrinovics et al., 2025). The structured, factually rich and linguistically minimal qualities of the KGs can potentially decrease the risks of the Needle-in-a-Haystack problem and limitations of the context window size. Conditioning LLMs on KGs can also enable optimal output scrutiny and explainability by allowing the outputs to be traced back to explicit sources, making cross-checking less time-consuming than document-based RAG. Furthermore, KGs accompany each entity with rich metadata, but their optimal use in factual language modeling is still an open question.
@@
Although KG-RAG is rapidly gaining attention to improve the factuality in LLM, existing QA benchmark data sets (Zhao et al., 2023; Lin et al., 2021; Li et al., 2023; Wei et al., 2024; Rahman et al., 2024; Mickus et al., 2024; Ravi et al., 2024; Bang et al., 2025) on LLM hallucinations rely primarily on textual data for contextual information and provide no multilingual support. While the questions in these benchmark datasets are compiled from different sources, the answers for FELM (Zhao et al., 2023) HaluEval  (Li et al., 2023) Shroom2024  (Mickus et al., 2024) are LLM-generated and evaluated using LLM-as-a-judge or human annotation. For some datasets (Lin et al., 2021; Zhao et al., 2023; Wei et al., 2024), the answers are supported with external contextual information from textual resources such as webpages. Therefore, in this paper, we bridge these critical gaps by presenting a novel multilingual hallucination benchmark MultiHal, grounded on factual information from Wikidata (Vrandečić & Krötzsch, 2014) KG. MultiHal is based on a total of 7 common benchmarks that lack structured factual and multilingual coverage, namely Felm (Zhao et al., 2023), TruthfulQA(Lin et al., 2021) (TQA), HaluEval(Li et al., 2023), HaluBench (Ravi et al., 2024), SimpleQA (Wei et al., 2024), DefAn (Rahman et al., 2024), Shroom2024 (Mickus et al., 2024). We propose a data collection framework as illustrated in Figure 1, to aggregate over 31k unique questions from aforementioned datasets, enriching them by mining 140k KG paths and ensuring factual consistency by filtering using LLM-as-a-judge. To enable multilingual hallucination evaluation, our compiled dataset comprising questions, ground-truth answers and KG paths, is translated to Spanish, French, Italian, Portuguese and German. Therefore, our main contributions are as follows:
@@ 2 MultiHal
MultiHal builds upon a set of 7 previously established benchmarks by enriching them with factual information in the form of relevant paths from Wikidata. The choice of these benchmarks is motivated by their relevancy to factuality evaluation, yet they lack support for factual grounding of the answers, leveraging KG and LLM integration models, and multilingual evaluation. We summarize the basic dataset statistics in Table 1, for MultiHal a dataset schema description see Appendix A. These foundational benchmarks are all filtered for generative question-answering based on general/trivia domains. Furthermore, benchmarks such as Shroom2024 (Mickus et al., 2024), FELM (Zhao et al., 2023), HaluEval (Li et al., 2023), HaluBench (Ravi et al., 2024) are primarily oriented towards evaluating hallucination detection models consisting of both hallucinated and non-hallucinated data, therefore, the data is repurposed by filtering for rows labelled as non-hallucinated. We consider that each unique question-path pair as a data point. The count difference between data points and unique questions is due to multiple candidate paths per question. The overview of each of the processing stages in our dataset collection pipeline is illustrated in Figure 1. The following sections scope into the methodological details of each of the processing stages of the proposed dataset collection framework. Additionally we report on our computing processing times and CO2 emissions in Appendix K. Our original contributions are released under CC-BY-4.0 license terms.
@@
Shroom2024 Definition Modeling $C$-BY 346 160 1 170.86 73.37
Felm World knowledge $C$-BY-NC-SA-4.0 73 17 1 95.25 75.26
MultiHal (total) - $C$-BY-4.0 25,905 7095 48 106.27 70.98
@@ 2.1 Dataset Preprocessing
Considering that MultiHal builds upon established benchmarks, question deduplication is performed to avoid data leakage across the foundations. Deduplication is based on computing sentence embeddings using SentenceTransformers [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 (Reimers & Gurevych, 2019) and computing all possible pair-wise cosine similarity between the questions. The ground-truth answers and any present supplementary context of the pair of questions with a sentence similarity threshold above 0.99 are merged. Deduplication was exclusively skipped for DefAn-QSRanking subset due to a large amount of questions consisting of nearby years for corresponding university rankings, which led to a very high number of false positives among the data points.
Additionally, we discard data points where the ground-truth answers are phrases such as "I have no comment", which indicate refusal to answer, and we define them as refusal types. We compile a list of refusals consisting of a list of text patterns as described in C. Any rows with output columns that exactly match, case-insensitively, one of these refusal phrases are filtered out.
@@ 2.2 KG Path Mining
The overall idea is to mine relevant paths from Wikidata (Vrandečić & Krötzsch, 2014). The core semantic entities are extracted from a given question $\mathcal{Q}$ and its ground-truth answer $\mathcal{A}$, and afterward matched to Wikidata entities. The extracted entities in $\mathcal{Q}$ and $\mathcal{A}$ are used for querying Wikidata for existing paths.
Entity Matching from Text to Knowledge Graphs. The core entity extraction and matching from raw text is based on Falcon 2.0 (Sakor et al., 2020). Falcon 2.0 is an open-sourced framework which is also made available via an API [^2] https://labs.tib.eu/falcon/falcon2/ that we call to retrieve subjects from question $\mathcal{Q}$ and objects from answer $\mathcal{A}$. Given a text passage, Falcon 2.0 outputs a ranked list of entities as candidates in Wikidata, we use the Top-3 candidates. For increased redundancy, we use Falcon 2.0 to additionally return DBpedia entities, which we then map back to Wikidata using the query in Listing 1 in Appendix D. Additionally, foundational benchmarks such as FELM (Zhao et al., 2023), SimpleQA (Wei et al., 2024), TruthfulQA (Lin et al., 2021) contain supplementary context in the form of Wikipedia links which we map to Wikidata(Vrandečić & Krötzsch, 2014) entities using Wikipedia public API [^3] https://en.wikipedia.org/w/api.php?action=query&prop=pageprops&titles=\$WIKIPEDIA_ID&format=json. The Wikipedia-to-Wikidata retrieval is done by taking the page title embedded in the given Wikipedia link and replacing it with the \$WIKIPEDIA_ID placeholder.
@@
Knowledge Graph Querying. We query Wikidata in order to find existing paths between the extracted subject-object entities up to 2 hops. As additional pre-processing steps before querying, we remove circular subject-object, as well as create an inverted set of subject-object pairs to accommodate for the directionality of the Wikidata graph. Depending on the foundational benchmark, we create custom queries for the different answer types we encounter when merging all our foundational benchmarks. The answer type is is denoted by answer_type column in MultiHal, see the schema in Appendix A. In Appendix D, see Listing 2 for Wikidata entity query, Listing 3 for date-literal query and Listing 4 for numerical-literal query. The answer types, such as numericals and dates, are queried with value limitations for numerical and time-based properties in the final hop, as shown in Listings 4 and 3 to improve query speed. See Appendix E Listing 6 for a set of time-based properties and Listing 7 for numerical properties. For querying, we use the public Wikidata endpoint [^4] https://query.wikidata.org/bigdata/namespace/wdq/sparql, our path cut-off date is April 2025.
For decoding the Wikidata entity labels, we run a separate pass using the query in Appendix D Listing 5. When querying for labels, we discard any statements, entities or objects that cannot be directly mapped to natural language text labels.
@@ 2.3 KG Path Quality Evaluation: LLM as Judge
As a method for filtering out noisy KG paths and identifying high-quality paths, we employ a two-step LLM-as-a-judge methodology (Li et al., 2025; Yu et al., 2024): firstly, for questions with more than 10 candidate paths the top-10 paths selection is done to limit the total count; secondly, scoring each path individually to identify low and high quality paths. For both selection and scoring we use GPT-4o Mini similarly to Laskar et al. (2025); Arif et al. (2025). We further motivate the choice of GPT-4o-Mini in Section 4.1. For inference we use OpenRouter API [^5] https://openrouter.ai/openai/gpt-4o-mini with sampling temperature 0.1. The goal of the selection is to decrease the overall number of KG paths, resulting in a decrease from 140k to 25.9k paths.
Selection Step. We construct a prompt for selecting the 10 most relevant paths with respect to the question-answer, and available optional answer pairs. Selection is intended to be done without any particular ordering, see Listing 8 in Appendix F. The set of paths for each question is processed in two passes, and in both scenarios, the order of the paths is shuffled to avoid any ordering biases (Li et al., 2025). From the two passes, we consider only the overlapping paths from the two candidate sets as the final collection of paths. During the selection phase, LLM-generated outputs are validated by checking for exact matches against the set of paths corresponding to the question. Any generated paths that do not have an exact match are discarded to mitigate the risk of syntactic errors or hallucinations introduced by the LLM-as-a-judge as a method of quality control. This process is repeated up to three times or until a total of 10 valid paths are obtained. If, after three attempts, fewer than 10 valid paths are selected, the remaining slots are filled by randomly sampling from the original KG path pool for the corresponding question. The selection step is bypassed for questions that have 10 or fewer candidate paths.
Scoring Step. Once a set of candidate paths is established, we construct another prompt for rating their relevance with respect to the given question and answer, and we process each path individually, see Appendix F Listing 9 for the instructions. Scoring is done by determining the quality score on a scale of 1-5, where 1 indicates a path which is completely unrelated to the question and answer, and 5 indicates an explicit answer to the question. From our final benchmark we filter out all paths rated 1-3, which we deem as low-quality and leave only paths rated 4-5 as high-quality ones.